# Faithfulness and the Notion of Adversarial Sensitivity in NLP Explanations

### v1

@@ Abstract
Faithfulness is a critical metric to assess the reliability of explainable AI. In NLP, current methods for faithfulness evaluation are fraught with discrepancies and biases, often failing to capture the true reasoning of models. We introduce Adversarial Sensitivity as a novel approach to faithfulness evaluation, focusing on the explainer’s response when the model is under adversarial attack. Our method accounts for the faithfulness of explainers by capturing sensitivity to adversarial input changes. This work addresses significant limitations in existing evaluation techniques, and furthermore, quantifies faithfulness from a crucial yet underexplored paradigm.00footnotetext: Accepted as a Full Paper at EMNLP 2024 Workshop BlackBoxNLP; reviews can be found here: https://openreview.net/forum?id=FD25pUum9k&nesting=2&sort=date-desc
@@
Supriya Manna SRM University AP reachsmanna@gmail.com                        Niladri Settthanks: all correspondences to settniladri@gmail.com SRM University AP settniladri@gmail.com
@@ 2 Adversarial Sensitivity
Adversarial Sensitivity and Faithfulness: Given $x^{\prime}$ is an AE of x for f, if I is ‘faithful’ to f, then $W_{x,f}$ and $W_{x^{\prime},f}$ should be dissimilar. In our setup, we report the mean distance over all obtained pairs of ($x,x^{\prime}$). This is a necessary but not sufficient condition for faithfulness. Currently (at the time of writing this paper) there is no necessary and sufficient condition for faithfulness (Lyu et al., 2024). However, following the argument of Lyu et al. (Lyu et al., 2024), as these metrics are primarily (meta)heuristic based evaluations, accessing faithfulness with several necessary tests is much practical than attempting to formulate an exhaustive list of necessary and sufficient conditions and then evaluating against all of them. Adversarial Sensitivity is one of such necessary tests to evaluate the faithfulness of explainers.
@@ 4.1 Datasets and Models
Model Dataset Accuracy (%) Accuracy after A1 (%) Accuracy after A2 (%) Accuracy after A3 (%) DERT SST-2 91.5 8.42 21.61 99.62 AG News 93.1 26.71 68.91 91.46 Twitter 51.7 18.84 8.28 96.93 BERT SST-2 92.43 10.77 19 99.42 AG News 94.40 25 32 94.50 Twitter 54.32 23.58 12.61 95.29

### v2

@@ Abstract
Faithfulness is a critical metric to assess the reliability of explainable AI. In NLP, current methods for faithfulness evaluation are fraught with discrepancies and biases, often failing to capture the true reasoning of models. We introduce Adversarial Sensitivity as a novel approach to faithfulness evaluation, focusing on the explainer’s response when the model is under adversarial attack. Our method accounts for the faithfulness of explainers by capturing sensitivity to adversarial input changes. This work addresses significant limitations in existing evaluation techniques, and furthermore, quantifies faithfulness from a crucial yet underexplored paradigm.
@@
Supriya Manna and Niladri Settthanks: Corresponding author SRM University AP, India reachsmanna@gmail.com, settniladri@gmail.com
@@ 2 Adversarial Sensitivity
Adversarial Sensitivity and Faithfulness: Given $x^{\prime}$ is an AE of x for f, if I is ‘faithful’ to f, then $W_{x,f}$ and $W_{x^{\prime},f}$ should be dissimilar. In our setup, we report the mean distance over all obtained pairs of ($x,x^{\prime}$). This is a necessary but not sufficient condition for faithfulness. Currently (at the time of writing this paper) there is no necessary and sufficient condition for faithfulness (Lyu et al., 2024). However, following the argument of Lyu et al. (Lyu et al., 2024), as these metrics are primarily (meta)heuristic based evaluations, accessing faithfulness with several necessary tests is much more practical than attempting to formulate an exhaustive list of necessary and sufficient conditions and then evaluating against all of them. Adversarial Sensitivity is one of such necessary tests to evaluate the faithfulness of explainers.
@@ 4.1 Datasets and Models
Model Dataset Accuracy (%) Accuracy after A1 (%) Accuracy after A2 (%) Accuracy after A3 (%) DERT SST-2 91.50 8.42 21.61 99.62 AG News 93.10 26.71 68.91 91.46 Twitter 51.7 18.84 8.28 96.93 BERT SST-2 92.43 10.77 19.00 99.42 AG News 94.40 25.00 32.00 94.50 Twitter 54.32 23.58 12.61 95.29